A typical reaction between an antacid and the hydrochloric acid in gastric juice is

NaHCO3(s) + HCl(aq)<=>NaCl(aq) + H2O(l) + CO2(g)

Calculate the volume (in L) of CO2 gas generated from 0.350 g of NaHCO3 and excess gastric juice at 1.00 atm and 37.0°C.
1
Expert's answer
2020-12-01T09:03:42-0500

N(NaHCO3)=N(CO2)=0.35/84=0.00417 mole

PV=NRT

V=NRT/P=0.00417*8.314*298/101,325=0.102 liter

V(CO2)=V(HCl)=0.102 liter